3. A car moving with constant acceleration covers the distance between two points 60 m apart in 6 s. Its velocity as it passes the second point is 15 m/s.
a) What is its velocity at the first point?
b) What is the acceleration?
1
Expert's answer
2020-12-08T10:53:29-0500

"s=\frac{v^2-v_0^2}{2a}" and "v=v_0+at" "\to" "v_0=v-at"



"s=\frac{v^2-(v_0-at)^2}{2a}\to a=\frac{2vt-2s}{t^2}=\frac{2\cdot 15\cdot 6-2\cdot60}{6^2}=1.67(m/s^2)"


"v_0=v-at=15-1.67\cdot6=5(m/s)"


Answer. "v_0=5(m\/s); a=1.67(m\/s^2)"